# Expander estimates for cubes

### v1

@@ Abstract.
00footnotetext: Keywords: sumsets, exponential density, differencing, exponential sums.00footnotetext: MSC(2020): 11B13, 11P5500footnotetext: The first author was supported by Deutsche Forschungsgemeinschaft (DFG) project numbers 255083470 and 462335009. The second author was supported by DFG project number 255083470 and a Leverhulme Early Career Fellowship, and received support from the Max Planck Institute for Mathematics and from the European Reseach Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant ID 648329). For the purpose of open access, the author has applied a Creative Commons Attribution (CC-BY) licence to any Author Accepted Manuscript version arising from this submission.
@@ 1. Introduction
The definition of $\delta$ implies $A\ll N^{\delta+\varepsilon}$, whence provided that $\delta\leq 1/k$, the right hand side of (1.4) is $O(N^{1/k+3\varepsilon}A)$. By (1.3) we finally see that
$\sum_{{\scriptstyle n\leq 2N}\atop{\scriptstyle r(n,N)\geq 1}}1\gg AN^{1/k-3\varepsilon}\gg N^{\delta+1/k-4\varepsilon}.$
@@
in integers $x,y,a,b$ with $1\leq x,y\leq N^{1/k}$ and $a,b\in\mathscr{A}$, $a,b\leq N$. The number of such solutions with $a=b$ is $[N^{1/k}]A$. Meanwhile, when $a\neq b$, a divisor function estimate shows that the number of solutions of (1.5) in $x,y$ is $O(N^{\varepsilon})$, and the number of choices for $a$, $b$ is $A^{2}$. This establishes (1.4) at once.
@@
If true, one may feed this bound to (1.3) to infer that (1.2) holds in general. This, our second, heuristical argument in support of (1.2), gives us reason to name (1.2) the expander conjecture for $k$-th powers. Despite its simple proof for $k=2$, and for the limited range $\delta\leq 1/k$ for larger $k$, no other case of our conjecture is currently known. The best lower bounds for $\delta_{k}$ are due to Davenport [4], dating back to 1942 (see also [12, Theorem 6.2]). In the cubic case $k=3$, he showed that if $\frac{1}{3}\leq\delta<1$, then one has
@@
Davenport used his lower bounds for $\delta_{k}$ in his work on Waring’s problem. With such applications in mind, various methods have been developed that improve on Davenport’s estimates for specific sets $\mathscr{A}$. In Waring’s problem, one may explore the homogeneity of sums of $k$-th powers. Davenport [4, Theorem 4] proposed one such method that was later developed by Vaughan [10, 11]. Early experiments in the applicability of the circle method are due to Vaughan [9] and Thanigasalam [7]. In Vaughan’s work it was important that the input set $\mathscr{A}$ stems from a polynomial, or at the very least, that $\mathscr{A}$ is somewhat regularly distributed over arithmetic progressions. He showed that sums of three positive cubes have exponential density at least $\frac{8}{9}$. We take $\mathscr{C}=\{x^{3}+y^{3}:\,x,y\in\mathbb{N}\}$, and note that $\delta(\mathscr{C})=\frac{2}{3}$. Our Theorem 1.1 yields $\delta_{3}(\mathscr{C})\geq\frac{8}{9}$ and recovers Vaughan’s result [^1] For the best currently known lower bound for $\delta_{3}(\mathscr{C})$ see [14]. This is not a coincidence. In fact, our Theorem 1.1 results from an attempt to trim Vaughan’s argument to apply to general input sets.
The simple argument that gave (1.2) is rather flexible. We modify it following Davenport [4] and Vaughan [9]. Choose $P=N^{2/5}$ and then examine the number of solutions of (1.5) with $k=3$ and $P<x,y\leq 2P$. Note that implicit in the equation and the size constraints is the condition $|x-y|\leq\surd P$. Like Vaughan, we apply the circle method to the equation (1.5), bring in $h=y-x$ as a new variable, and are then led to study the exponential sum
@@
[…]
@@
$F_{w}(\alpha)\ll HP(\log P)\Xi(\alpha)+P^{1+\varepsilon}.$
With Theorem 1.2 in hand, the proof of its corollary, named here Theorem 1.3, is so easy that we give it at once. In fact, by (1.9), we see that whenever $\alpha\in\mathfrak{M}\setminus\mathfrak{N}$, then $\Upsilon_{w}(\alpha)\ll P^{-1/2}$ holds uniformly in $w$. This shows that one has $F_{w}(\alpha)\ll P^{1+\varepsilon}$ for these $\alpha$. Further, when $\alpha\in\mathfrak{N}(q,a)$ for some coprime pair $a,q$ with $q\leq Q$ and $q$ has a cubic factor, then in particular, there is a prime $p$ with $p^{3}\mid q$. Now $p\leq q^{1/3}\leq P^{1/4}$, and hence $\kappa_{\varpi}(q)=0$. Consequently, in this case we again have $F_{\varpi}(\alpha)\ll P^{1+\varepsilon}$. This leaves the case where $q$ is cube-free. Then the inequality $\kappa_{w}(q)\leq 2^{\omega(q)}q^{-1/2}$ is immediate from the definition of $\kappa_{w}$, and the conclusion of Theorem 1.3 follows from Theorem 1.2.
@@
In this paper, we apply the following convention concerning the letter $\varepsilon$. Whenever $\varepsilon$ occurs in a statement, it is asserted that the statement is true for any positive real value assigned to $\varepsilon$. Constants implicit in the use of Vinogradov’s and Landau’s familiar symbols may depend on the value of $\varepsilon$. We abbreviate $\exp(2\pi\mathrm{i}\alpha)$ to $e(\alpha)$. The number of divisors of the natural number $q$ is denoted $\tau(q)$, and $\phi(q)$ is Euler’s totient.
@@ 3. Major arcs: first application of Poisson summation
Note here that for $d>H$ one has $G(\beta,P/d,H/d)=0$ so that implicitly the sum over $d$ is constrained to $d\leq H$. We shall invoke this comment frequently and silently. We require all our estimates uniformly relative to the auxiliary natural number $w$. Note here that primes $p$ with $p>H$ are coprime to all $h\leq H$. Hence, there is no loss of generality in assuming henceforth that $w$ is a square-free number free of prime factors exceeding $H$.
@@
$D=D(\alpha,w)=\{d\in{\mathbb{N}}:d\mid w,\,d\leq H,\,d|\beta|\leq(48P^{2})^{-1}\}$ (3.8)
@@ 4. Major arcs: second application of Poisson summation
In the interest of brevity, we write $U(q,a,0)=U(q,a)$ for the cubic Gauss sum.
@@
that holds for coprime natural numbers $q_{1},q_{2}$ and any integers $b,c$. This is proved en passant on p. 38 of Vaughan [12], subject to the condition that $(c,q_{1}q_{2})=1$. However, an inspection of the underlying argument shows that this extra condition is not required,
@@
$\sum_{h\leq H/d}\!e(\Phi)=\sum_{k=1}^{q}e\Big{(}\frac{ad^{3}(k^{3}+3zk^{2}+3z^{2}k)}{q}\Big{)}\sum_{{\scriptstyle h\leq H/d}\atop{\scriptstyle h\equiv k\bmod k}}\!e\big{(}\beta(d^{3}h^{3}+3ud^{2}h^{2}+3u^{2}dh)\big{)}.$
@@
$F^{*}(\alpha)=\sum_{d\in D}\frac{\mu(d)}{q^{2}d}\sum_{-q/2<b\leq q/2}T(q,ad^{3},b)\int_{P}^{2P}\sum_{h\leq H/d}e(\beta\Psi(u,dh)-bh/q)\,{\mathrm{d}}u$ (4.8)
@@
$\int_{P}^{2P}\!\!\!\int_{0}^{H/d}e(\beta\Psi(u,d\eta))\,{\mathrm{d}}\eta\,{\mathrm{d}}u=\frac{1}{d}\int_{0}^{H}\!\!\!\int_{P}^{2P}e(\beta\Psi(u,d\eta))\,{\mathrm{d}}u\,{\mathrm{d}}\eta=\frac{K(\beta)}{d}.$
@@
$\psi(\eta)=\beta\Psi(u,d\eta)-c\eta/q=\beta(d^{3}\eta^{2}+3ud^{2}\eta^{2}+3u^{2}d\eta)-c\eta/q$
defines a twice differentialable function on $[0,H/d]$, and one has
@@
uniformly for all $u\in[P,2P]$. Let $C_{d}^{+}$ be the set of all integers $c\neq 0$ with $-\frac{3}{2}q<c\leq\frac{3}{2}q$ where (4.13) holds for all $u\in[P,2P]$ and all $\eta\in[0,H/d]$. Then, by Lemma 4.2, the terms with $c\in C_{d}^{+}$ contribute to (4.12) an amount not exceeding
@@
Let $C_{d}^{-}$ be the set of all non-zero integers $c$ with $-\frac{3}{2}q<c\leq\frac{3}{2}q$ that are not in $C_{d}^{+}$. Recalling (4.13), we see that for each $c\in C_{d}^{-}$, there are real numbers $\eta,u$ with $0\leq\eta\leq H/d$, $P\leq u\leq 2P$ and
@@
is in the interval $[\frac{4}{5},\frac{6}{5}]$. We conclude that for large $P$ we have
@@
This shows that $C_{d}^{-}$ is contained in the interval $[1,16qd\beta P^{2}]$. Recall that $d\in D$, whence $d\beta\leq 1/(48P^{2})$, so that we have gained new information here. The total contribution of the sets $C_{d}^{-}$ to (4.12) now is
@@
In this analysis, we supposed that $\beta>0$. If $\beta<0$, then we may change $\beta$ and $c$ in (4.15) to $-\beta$ and $-c$ to realize that the same argument applies. By (4.14) and (4.16) we now see that terms with $c\neq 0$ in (4.12) contribute at most $O(P^{1+\varepsilon})$, as required to complete the proof of Lemma 4.1.
@@ 5. Major arcs: the local factors
We begin by collecing a number of well known results on cubic Gauss sums. First we note that, as an instance of (4.3), when $q_{1}$, $q_{2}$ are coprime natural numbers and $c$ is an integer, then
@@
Given $q$ and the square-free number $w$, we put $w_{0}=(q,w)$ and $w=w_{0}w_{1}$. Then $(w_{0},w_{1})=1$, so that each $d$ dividing $w$ factors uniquely as $d=d_{0}d_{1}$ with $d_{0}\mid w_{0}$, $d_{1}\mid w_{1}$. Furthermore, let $q_{1}$ denote the largest divisor of $q$ coprime to $w$, and put $q=q_{0}q_{1}$. Then the prime factors of $q_{0}$ divide $w_{0}$. In particular, $(q_{0},q_{1})=1$. From (5.2) we now infer
$\mathscr{S}(a/q,w)=\sum_{d_{0}\mid w_{0}}\sum_{d_{1}\mid w_{1}}\frac{\mu(d_{0})\mu(d_{1})}{q_{0}^{2}q_{1}^{2}}|U(q_{0},aq_{1}^{2}d^{3})U(q_{1},aq_{0}^{2}d^{3})|^{2}.$
@@
$W(r,b)=\sum_{d\mid r}\mu(d)\sum_{{\scriptstyle x,y=1}\atop{\scriptstyle d\mid x,\,d\mid y}}^{r}e\Big{(}\frac{b}{r}(x^{3}-y^{3})\Big{)}=\sum_{d\mid r}\mu(d)|U(r/d,bd^{3})|^{2}.$
By periodicity, for any $c\in{\mathbb{Z}}$, we have $U(r/d,c)=d^{-1}U(r,c)$, and then
@@
The sum $W(r,b)$ is quasi-multiplicative in a manner similar to (5.5). In fact, when $r=r_{1}r_{2}$ with $(r_{1},r_{2})=1$, the numbers $r_{1}x_{2}+r_{2}x_{1}$ with $1\leq x_{j}\leq r_{j}$ $(j=1,2)$ meet each residue class modulo $r_{1}r_{2}$ exacly once. We now substitute
@@
in (5.9). The condition $(x,y,r)=1$ is equivalent with $(x_{1},y_{1},r_{1})=(x_{2},y_{2},r_{2})=1$, and one observes that

### v2

@@ Abstract.
00footnotetext: Keywords: sumsets, exponential density, differencing, exponential sums.00footnotetext: MSC(2020): 11B13, 11P5500footnotetext: The first author was supported by Deutsche Forschungsgemeinschaft (DFG) project numbers 255083470 and 462335009. The second author was supported by DFG project number 255083470 and a Leverhulme Early Career Fellowship, and received support from the Max Planck Institute for Mathematics and from the European Reseach Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant ID 648329). For the purpose of open access, the author has applied a Creative Commons Attribution ($C$-BY) licence to any Author Accepted Manuscript version arising from this submission.
@@ 1. Introduction
The definition of $\delta$ implies $A\ll N^{\delta+\varepsilon}$, whence provided that $\delta\leq 1/k$, the right hand side of (1.4) is $O(N^{1/k+2\varepsilon}A)$. By (1.3) we finally see that
$\sum_{{\scriptstyle n\leq 2N}\atop{\scriptstyle r(n,N)\geq 1}}1\gg AN^{1/k-2\varepsilon}\gg N^{\delta+1/k-3\varepsilon}.$
@@
in integers $x,y,a,b$ with $1\leq x,y\leq N^{1/k}$ and $a,b\in\mathscr{A}$, $a,b\leq N$. The number of such solutions with $a=b$ is $N^{1/k}]A$. Meanwhile, when $a\neq b$, a divisor function estimate shows that the number of solutions of (1.5) in $x,y$ is $O(N^{\varepsilon})$, and the number of choices for $a$, $b$ is $A^{2}$. This establishes (1.4) at once.
@@
If true, one may feed this bound to (1.3) to infer that (1.2) holds in general. This second heuristical argument in support of (1.2) gives us reason to name (1.2) the expander conjecture for $k$-th powers. Despite its simple proof for $k=2$, and for the limited range $\delta\leq 1/k$ for larger $k$, no other case of our conjecture is currently known. The best lower bounds for $\delta_{k}$ are due to Davenport [4], dating back to 1942 (see also [12, Theorem 6.2]). In the cubic case $k=3$, he showed that if $\frac{1}{3}\leq\delta<1$, then one has
@@
Davenport used his lower bounds for $\delta_{k}$ in his work on Waring’s problem. With such applications in mind, various methods have been developed that improve on Davenport’s estimates for specific sets $\mathscr{A}$. In Waring’s problem, one may explore the homogeneity of sums of $k$-th powers. Davenport [4, Theorem 4] proposed one such method that was later developed by Vaughan [10, 11]. Early experiments in the applicability of the circle method are due to Vaughan [9] and Thanigasalam [7]. In Vaughan’s work it was important that the input set $\mathscr{A}$ stems from a polynomial, or at the very least, that $\mathscr{A}$ is somewhat regularly distributed over arithmetic progressions. He showed that sums of three positive cubes have exponential density at least $\frac{8}{9}$. We take $\mathscr{C}=\{x^{3}+y^{3}:\,x,y\in\mathbb{N}\}$, and note that $\delta(\mathscr{C})=\frac{2}{3}$. Our Theorem 1.1 yields $\delta_{3}(\mathscr{C})\geq\frac{8}{9}$ and recovers Vaughan’s result [^1] For the best currently known lower bound for $\delta_{3}(\mathscr{C})$ see [14].. This is not a coincidence. In fact, our Theorem 1.1 results from an attempt to trim Vaughan’s argument to apply to general input sets.
The simple argument that gave (1.2) is rather flexible. We modify it following Davenport [4] and Vaughan [9]. Choose $P=N^{2/5}$ and then examine the number of solutions of (1.5) with $k=3$ and $P<x,y\leq 2P$. Note that implicit in the equation and the size constraints is the condition $|x-y|\leq\surd P$. Like Vaughan, we apply the circle method to the equation (1.5), bring in h=yxh=y-x as a new variable, and are then led to study the exponential sum
@@
As it turns out, the condition $(x,y,w)=1$ eliminates unwanted peaks of (1.7) at $a/q$ as long as the cubic factors of $q$ are solely composed of primes dividing $w$. In order to describe this precisely, we need to introduce more notation. The estimate for $F_{w}$ features prominently a certain multiplicative function $\kappa_{w}$ that we define on powers of a prime $p$ with $p\nmid w$ by
@@
$F_{\varpi}(\alpha)\ll HP(\log P)\Xi(\alpha)+P^{1+\varepsilon}.$
With Theorem 1.2 in hand, the proof of its corollary, named here Theorem 1.3, is so easy that we give it at once. In fact, by (1.9), we see that whenever $\alpha\in\mathfrak{M}\setminus\mathfrak{N}$, then $\Upsilon_{w}(\alpha)\ll P^{-1/2}$ holds uniformly in $w$. This shows that one has $F_{w}(\alpha)\ll P^{1+\varepsilon}$ for these $\alpha$. Further, when $\alpha\in\mathfrak{N}(q,a)$ for some coprime pair $a,q$ with $q\leq Q$ and $q$ has a cubic factor, then in particular, there is a prime $p$ with $p^{3}\mid q$. Now $p\leq q^{1/3}\leq P^{1/4}$, and hence $\kappa_{\varpi}(q)=0$, where $\varpi$ is as in Theorem 1.3. Consequently, in this case we again have $F_{\varpi}(\alpha)\ll P^{1+\varepsilon}$. This leaves the case where $q$ is cube-free. Then the inequality $\kappa_{\varpi}(q)\leq 2^{\omega(q)}q^{-1/2}$ is immediate from the definition of $\kappa_{w}$, and the conclusion of Theorem 1.3 follows from Theorem 1.2.
@@
In this paper, we apply the following convention concerning the letter $\varepsilon$. Whenever $\varepsilon$ occurs in a statement, it is asserted that the statement is true for any positive real value assigned to $\varepsilon$. Constants implicit in the use of Vinogradov’s and Landau’s familiar symbols may depend on the value of $\varepsilon$. We abbreviate $\exp(2\pi\mathrm{i}\alpha)$ to $e(\alpha)$. The number of divisors of the natural number $q$ is denoted $\tau(q)$, and $\phi(q)$ is Euler’s totient function. We reserve the letter $d$ to denote a square-free number; we may also use $d_{0},d^{\prime}$ and so forth. As our main objects of study, $F_{w}$ and $\kappa_{w}$, depend on $w$ only on through its square-free part, we will invariably view $w$ through the lens of its square-free divisors $d$.
An anonymous referee proposed a substantial number of minor alterations to an earlier version of this text. These were evidently thought through with care, and we feel the result is quite a bit more readable. We gratefully acknowledge this invaluable help.
@@ 3. Major arcs: first application of Poisson summation
Note here that for $d>H$ one has $G(\beta,P/d,H/d)=0$ so that implicitly the sum over $d$ is constrained to $d\leq H$. We shall invoke this comment frequently and silently. We require all our estimates uniformly relative to the auxiliary natural number $w$. Note here that primes $p$ with $p>H$ are coprime to all $h\leq H$. Hence, there is no loss of generality in assuming henceforth that $w$ is a number free of prime factors exceeding $H$.
@@
$D=D(\alpha,w)=\{d\in{\mathbb{N}}:\mu(d)^{2}=1,\,d\mid w,\,d\leq H,\,d|\beta|\leq(48P^{2})^{-1}\}$ (3.8)
@@ 4. Major arcs: second application of Poisson summation
In the interest of brevity, we write $U(q,a,0)=U(q,a)$ for the cubic Gauss sum. We continue to denote by $D$ the particular set of square-free numbers defined in (3.8).
@@
that holds for coprime natural numbers $q_{1},q_{2}$ and any integers $b,c$. This is proved en passant on p. 38 of Vaughan [12], subject to the condition that $(c,q_{1}q_{2})=1$. However, an inspection of the underlying argument shows that this extra condition is not required.
@@
$\sum_{h\leq H/d}\!e(\Phi)=\sum_{k=1}^{q}e\Big(\frac{ad^{3}(k^{3}+3zk^{2}+3z^{2}k)}{q}\Big)\sum_{{\scriptstyle h\leq H/d}\atop{\scriptstyle h\equiv k\bmod q}}\!e\big(\beta(d^{3}h^{3}+3ud^{2}h^{2}+3u^{2}dh)\big).$
@@
$F_{w}^{*}(\alpha)=\sum_{d\in D}\frac{\mu(d)}{q^{2}d}\sum_{-q/2<b\leq q/2}T(q,ad^{3},b)\int_{P}^{2P}\sum_{h\leq H/d}e(\beta\Psi(u,dh)-bh/q)\,{\mathrm{d}}u$ (4.8)
@@
$\int_{P}^{2P}\!\!\!\int_{0}^{H/d}e(\beta\Psi(u,d\eta))\,{\mathrm{d}}\eta\,{\mathrm{d}}u=\frac{1}{d}\int_{0}^{H}\!\!\!\int_{P}^{2P}e(\beta\Psi(u,\eta))\,{\mathrm{d}}u\,{\mathrm{d}}\eta=\frac{K(\beta)}{d}.$
@@
$\psi(\eta)=\beta\Psi(u,d\eta)-c\eta/q=\beta(d^{3}\eta^{3}+3ud^{2}\eta^{2}+3u^{2}d\eta)-c\eta/q$
defines a twice differentiable function on $[0,H/d]$, and one has
@@
uniformly for all $u\in[P,2P]$. Let $C_{d}^{+}$ be the set of all integers $c\neq 0$ with $\frac{3}{2}q<c\leq\frac{3}{2}q$ where (4.13) holds for all $u\in[P,2P]$ and all $\eta\in[0,H/d]$. Then, by Lemma 4.2, the terms with $c\in C_{d}^{+}$ contribute to (4.12) an amount not exceeding
@@
Let $C_{d}^{-}$ be the set of all non-zero integers $c$ with $\frac{3}{2}q<c\leq\frac{3}{2}q$ that are not in $C_{d}^{+}$. Recalling (4.13), we see that for each $c\in C_{d}^{-}$, there are real numbers $\eta,u$ with $0\leq\eta\leq H/d$, $P\leq u\leq 2P$ and
@@
is in the interval $\frac{4}{5},\frac{6}{5}]$. We conclude that for large $P$ we have
@@
This shows that $C_{d}^{-}$ is contained in the interval $1,16qd\beta P^{2}]$. Recall that $d\in D$, whence $d\beta\leq 1/(48P^{2})$, so that we have gained new information here. The total contribution of the sets $C_{d}^{-}$ to (4.12) now is
@@
In this analysis, we supposed that $\beta>0$. If $\beta<0$, then we may change $\beta$ and $c$ in (4.15) to $\beta$ and c-c to realize that the same argument applies. By (4.14) and (4.16) we now see that terms with $c\neq 0$ in (4.12) contribute at most $O(P^{1+\varepsilon})$, as required to complete the proof of Lemma 4.1.
@@ 5. Major arcs: the local factors
We begin by collecting a number of well known results on cubic Gauss sums. First we note that, as an instance of (4.3), when $q_{1}$, $q_{2}$ are coprime natural numbers and $c$ is an integer, then
@@
Given $q$ and $w$, we put $w_{1}$ for the largest divisor of $w$ with $(q,w_{1})=1$, and write $w=w_{0}w_{1}$. Then $(w_{0},w_{1})=1$, so that each $d$ dividing $w$ factors uniquely as $d=d_{0}d_{1}$ with $d_{0}\mid w_{0}$, $d_{1}\mid w_{1}$. Furthermore, let $q_{1}$ denote the largest divisor of $q$ coprime to $w$, and put $q=q_{0}q_{1}$. Then the prime factors of $q_{0}$ divide $w_{0}$. In particular, $(q_{0},q_{1})=1$. From (5.2) we now infer
$\mathscr{S}(a/q,w)=\sum_{d_{0}\mid w_{0}}\sum_{d_{1}\mid w_{1}}\frac{\mu(d_{0})\mu(d_{1})}{q_{0}^{2}q_{1}^{2}d_{0}^{2}d_{1}^{2}}|U(q_{0},aq_{1}^{2}d^{3})U(q_{1},aq_{0}^{2}d^{3})|^{2}.$
@@
$W(r,b)=\sum_{d\mid r}\mu(d)\sum_{{\scriptstyle x,y=1}\atop{\scriptstyle d\mid x,\,d\mid y}}^{r}e\Big(\frac{b}{r}(x^{3}-y^{3})\Big)=\sum_{d\mid r}\mu(d)|U(r/d,bd^{2})|^{2}.$
By periodicity, for any $c\in{\mathbb{Z}}$, we have $U(r/d,c)=d^{-1}U(r,cd)$, and then
@@
The sum $W(r,b)$ is quasi-multiplicative in a manner similar to (5.5). In fact, when $r=r_{1}r_{2}$ with $(r_{1},r_{2})=1$, the numbers $r_{1}x_{2}+r_{2}x_{1}$ with $1\leq x_{j}\leq r_{j}$ $(j=1,2)$ meet each residue class modulo $r_{1}r_{2}$ exactly once. We now substitute
@@
in (5.9). The condition $(x,y,r)=1$ is equivalent to $(x_{1},y_{1},r_{1})=(x_{2},y_{2},r_{2})=1$, and one observes that